# Adapting Differentially Private Synthetic Data to Relational Databases

### v1

@@ 3 Main Results
We present our main algorithm (Algorithm 1) for generating privacy-preserving synthetic relational databases. This algorithm can be combined with any existing single-source DP synthetic data mechanisms, adapting them to the relational database context by establishing relationships among individual synthetic tables. Specifically, it learns a bi-adjacency matrix by minimizing approximation errors in cross-table $k$-way marginal queries compared to the original data. Given the inherent large size of the query class, our algorithm employs an iterative approach to identify the queries with the highest approximation errors and refine the bi-adjacency matrix to reduce these errors. Our algorithm is designed for efficiency and scalability, making it well-suited for high-dimensional data. It also ensures referential integrity in the generated synthetic relational data. We end this section by establishing rigorous utility and DP guarantees for our algorithm.
We first create individual synthetic tables $\mathcal{D}_{i}^{\text{syn}}$ by applying any DP mechanisms via black-box access. Our main technical contribution lies in establishing connections among these synthetic tables by constructing a bi-adjacency matrix $\bm{B}$ that links them. We learn this matrix by aligning the cross-table $k$-way marginal queries between the synthetic and original databases. The key observation is that each of these marginal queries can be written as a (fractional) linear function of the bi-adjacency matrix. We formalize this observation in the following lemma.
Lemma 1.
For a relational database $(\mathcal{D}_{1},\mathcal{D}_{2},\bm{B})$ and any $k$-way cross-table query $q_{(\mathcal{S}_{1},\mathcal{S}_{2}),(\bm{y}_{1},\bm{y}_{2})}$, let $\bm{1}_{(\mathcal{S}_{1},\bm{y}_{1})}\in\{0,1\}^{n_{1}}$ denote an indicator vector whose i-th element equals $1$ iff the $i$-th record in $\mathcal{D}_{1}$ satisfies $x_{i,j}=y_{1,j}$ for all $j\in\mathcal{S}_{1}$. Let $\bm{q}_{(\mathcal{S}_{1},\mathcal{S}_{2}),(\bm{y}_{1},\bm{y}_{2})}=\operatorname{vec}(\bm{1}_{(\mathcal{S}_{1},\bm{y}_{1})}\otimes\bm{1}_{(\mathcal{S}_{2},\bm{y}_{2})})$. Then we have:
@@
Based on the above lemma, we write the cross-table query and its corresponding vector $\bm{q}_{(\mathcal{S}_{1},\mathcal{S}_{2}),(\bm{y}_{1},\bm{y}_{2})}$ interchangeably. We learn the bi-adjacency matrix of the synthetic database by solving the following optimization:
$\min_{\begin{subarray}{c}\bm{b}^{\text{syn}}\in\{0,1\}^{n^{\text{syn}}_{1}\cdot n^{\text{syn}}_{2}}\\ \bm{1}^{T}\bm{b}^{\text{syn}}=m^{\text{syn}}\end{subarray}}\max_{\bm{q}\in\mathcal{Q}_{\text{cross},k}}\left|\frac{1}{m^{\text{syn}}}\bm{q}^{T}\bm{b}^{\text{syn}}-a\right|.$ (4)
where $\bm{b}^{\text{syn}}=\operatorname{vec}(\bm{B}^{\text{syn}})$ and $a$ is the query value computed from the original real data. Since the minimization over $\bm{b}^{\text{syn}}$ in (4) is a combinatorial optimization, which is inherently challenging to solve, we solve a relaxed problem by allowing $\bm{b}^{\text{syn}}\in[0,1]^{n^{\text{syn}}_{1}\cdot n^{\text{syn}}_{2}}$ and use a randomized rounding algorithm to convert the obtained values back into integer values.
We present an iterative algorithm for solving the above optimization with DP guarantees. At each iteration, we apply the exponential mechanism to identify cross-table $k$-way workloads whose corresponding queries have the highest approximation errors between synthetic and real databases. Then we add isotropic Gaussian noise to their query values and update the (vectorized) bi-adjacency matrix $\bm{b}^{\text{syn}}$ to reduce these approximation errors. Specifically, at iteration $t$, we stack all the queries reported so far into a matrix $\hat{\bm{Q}}$ and let their noisy answers be a vector $\hat{\bm{a}}$. To update $\bm{b}^{\text{syn}}$, we use a mirror descent update with negative entropy mirror [8] to solve the following bounded-variable least-squares problem [42, 7] (see Appendix B for details):
@@
Suppose $\bm{b}^{\text{syn}}$ is the output from the preceding procedure, which we reshape into a matrix. Next, we introduce a randomized rounding algorithm to convert $\bm{b}^{\text{syn}}$ into $\{0,1\}^{n^{\text{syn}}_{1}\times n^{\text{syn}}_{2}}$ for defining the relationship between synthetic tables. Given that $\bm{b}^{\text{syn}}/m^{\text{syn}}$ forms a valid probability distribution over the domain $[n^{\text{syn}}_{1}]\times[n^{\text{syn}}_{2}]$, we sample random values from this distribution, rejecting those that have already been selected. We iterate this process until we obtain $m^{\text{syn}}$ distinct random values. These values correspond to the indices of the bi-adjacency matrix $\bm{B}^{\text{syn}}$ whose value is $1$.
We outline the above procedure in Algorithm 1 and provide additional details, including our privacy budget tracking, hyper-parameters, the updating rule for mirror descent, and discretization strategy, in Appendix B. We explore strategies for ensuring the scalability of our algorithm for synthesizing large-scale relational databases in Appendix E. We show how our above procedure can be adapted if the original database has a one-to-many relationship in Appendix D. We remark that iterative algorithms are a standard technique for query releasing in the DP literature [see 17, 18, 27, 1, 32, for examples in synthetic data/graph generation] and we extend its applications to establish relationships between synthetic tables. We end this section by establishing DP and utility guarantees for our algorithm.
@@
Input: private relational database $\mathcal{B}=(\mathcal{D}_{1},\mathcal{D}_{2},\bm{B})$; privacy budget $(\varepsilon_{\text{rel}},\delta_{\text{rel}})$; individual synthetic tables $\mathcal{D}_{1}^{\text{syn}}$, $\mathcal{D}_{2}^{\text{syn}}$; workloads per iteration $K$; number of iterations $T$; number of relationship in synthetic tables: $m^{\text{syn}}$
@@
     Append the queries from the new workloads to $\hat{\bm{Q}}$.
     Add noise to the new query answers and append them to $\hat{\bm{a}}$.
     Solve $\bm{b}^{\text{syn}}={\arg\min}_{\begin{subarray}{c}\bm{b}\in[0,1]^{n_{1}^{\text{syn}}\cdot n_{2}^{\text{syn}}}\\ \bm{1}^{T}\bm{b}=m^{\text{syn}}\end{subarray}}\lVert\frac{1}{m^{\text{syn}}}\hat{\bm{Q}}\bm{b}-\hat{\bm{a}}\rVert_{2}^{2}$
     $\bm{B}^{\text{syn}}=\operatorname{RandomizedRounding}(\bm{b}^{\text{syn}}.\text{reshape}(n_{1}^{\text{syn}},n_{2}^{\text{syn}}))$

### v2

@@ 3 Main Results
We present our main algorithm (Algorithm 1) for generating privacy-preserving synthetic relational databases. This algorithm can be combined with any existing DP synthetic data generation mechanisms, adapting them to produce relational databases by establishing relationships among individual synthetic tables. Specifically, it learns a bi-adjacency matrix by minimizing approximation errors in terms of cross-table $k$-way marginal queries compared to the original data. Given the inherent large size of the query class, our algorithm employs an iterative approach to identify the queries with the highest approximation errors and refine the bi-adjacency matrix to reduce these errors. Notably, our algorithm is designed for efficiency and scalability, making it well-suited for high-dimensional data. It also ensures referential integrity in the generated synthetic relational data. We end this section by providing an overview of the rest of the paper, which elaborates on the technical details for our algorithm.
Our key observation is that cross-table marginal queries can be written as a (fractional) linear function of the bi-adjacency matrix. This formulation enables learning the bi-adjacency matrix for a synthetic relational database through an optimization problem. We present this observation formally in the following lemma.
Lemma 1
For a relational database $(\mathcal{D}_{1},\mathcal{D}_{2},\bm{B})$ and any $k$-way cross-table query $q_{(\mathcal{S}_{1},\mathcal{S}_{2}),(\bm{y}_{1},\bm{y}_{2})}$, let $\bm{1}_{(\mathcal{S}_{1},\bm{y}_{1})}\in\{0,1\}^{n_{1}}$ denote an indicator vector whose i-th element equals $1$ iff the $i$-th record in $\mathcal{D}_{1}$ satisfies $x_{i,j}=y_{1,j}$ for all $j\in\mathcal{S}_{1}$ and $\bm{1}_{(\mathcal{S}_{2},\bm{y}_{2})}$ is defined in a similar manner. Let $\bm{q}_{(\mathcal{S}_{1},\mathcal{S}_{2}),(\bm{y}_{1},\bm{y}_{2})}=\operatorname{vec}(\bm{1}_{(\mathcal{S}_{1},\bm{y}_{1})}\cdot\bm{1}^{T}_{(\mathcal{S}_{2},\bm{y}_{2})})$. Then we have:
@@
Proof  See Appendix D.1.
Based on the above lemma, we write the cross-table query and its corresponding vector $\bm{q}_{(\mathcal{S}_{1},\mathcal{S}_{2}),(\bm{y}_{1},\bm{y}_{2})}$ interchangeably. We learn the bi-adjacency matrix for the synthetic database by solving the following optimization:
$\min_{\begin{subarray}{c}\bm{b}^{\text{syn}}\in\{0,1\}^{n^{\text{syn}}_{1}\cdot n^{\text{syn}}_{2}}\\ \bm{1}^{T}\bm{b}^{\text{syn}}=m^{\text{syn}}\end{subarray}}\max_{\bm{q}\in\mathcal{Q}_{\text{cross},k}}\left|\frac{1}{m^{\text{syn}}}\bm{q}^{T}\bm{b}^{\text{syn}}-a\right|,$ (4)
where $\bm{b}^{\text{syn}}=\operatorname{vec}(\bm{B}^{\text{syn}})$ and $a$ is the query value computed from the original real data. Since the minimization over $\bm{b}^{\text{syn}}$ in (4) is a combinatorial optimization, which is inherently challenging to solve, we solve a relaxed problem by allowing $\bm{b}^{\text{syn}}\in[0,1]^{n^{\text{syn}}_{1}\cdot n^{\text{syn}}_{2}}$ and use an unbiased sampling algorithm to convert the obtained values back into integer values.
We present an iterative algorithm for solving the above optimization with DP guarantees. At each iteration, we apply the exponential mechanism to identify cross-table $k$-way workloads whose corresponding queries have the highest approximation errors between synthetic and real databases. Then we add isotropic Gaussian noise to their query values and update the (vectorized) bi-adjacency matrix $\bm{b}^{\text{syn}}$ to reduce these approximation errors. Specifically, at iteration $t$, we stack all the queries reported so far into a matrix $\hat{\bm{Q}}$ and let their noisy answers be a vector $\hat{\bm{a}}$. To update $\bm{b}^{\text{syn}}$, we apply projected gradient descent to solve the following quadratic program (Section 4):
@@
Suppose $\bm{b}^{\text{syn}}$ is the output from the preceding procedure, which we reshape into a matrix $\bm{B}^{\text{syn}}$. Next, we introduce an unbiased sampling algorithm to convert each element of $\bm{B}^{\text{syn}}$ into $\{0,1\}$ for defining the relationship between synthetic tables (Section 5). To ensure the scalability of our algorithm, we propose a random slicing method that allows us to generate large-scale relational databases (Section 6). Finally, we analyze the utility of our algorithm in preserving $k$-way marginal queries (Section 7) and validate our algorithm through numerical experiments (Section 8).
We outline the above procedure in Algorithm 1 and provide additional details, including our privacy budget tracking and hyper-parameters, in Section B. Throughout the paper, we assume that the original relational database demonstrates a many-to-many relationship. In Appendix C, we apply our algorithm to accommodate one-to-many relationships, which, as shown, are significantly easier to manage. We remark that iterative algorithms are a standard technique for query releasing in the DP literature (see Gupta et al., 2012; Hardt et al., 2012; Liu et al., 2021b; Aydore et al., 2021; McKenna et al., 2022, for examples in synthetic data/graph generation) and we extend its applications to establish relationships between synthetic tables. We end this section by establishing a DP guarantee for our algorithm.
Theorem 1
Assume that each record in the original relational database has at most $d_{\max}$ relationships with records in the other table. We apply any DP mechanism with privacy budgets $(\epsilon_{1},\delta_{1})$ and $(\epsilon_{2},\delta_{2})$ to generate individual synthetic tables, respectively. Then, we use Algorithm  1 with privacy budget $(\epsilon_{\text{rel}},\delta_{\text{rel}})$ to build the relationships between them. The resulting synthetic relational database satisfies $(\epsilon_{1}+\epsilon_{2}+\epsilon_{\text{rel}},\delta_{1}+\delta_{2}+\delta_{\text{rel}})$-DP.
Proof  See Appendix D.2.
@@
Input: private relational database $\mathcal{B}=(\mathcal{D}_{1},\mathcal{D}_{2},\bm{B})$; privacy budget $(\epsilon_{\text{rel}},\delta_{\text{rel}})$; individual synthetic tables $\mathcal{D}_{1}^{\text{syn}}$, $\mathcal{D}_{2}^{\text{syn}}$; workloads per iteration $K$; number of iterations $T$; number of relationship in synthetic tables $m^{\text{syn}}$
@@
    Solve the relaxed optimization $\widetilde{\bm{b}}^{\text{syn}}={\arg\min}_{\begin{subarray}{c}\bm{b}\in[0,1]^{n_{1}^{\text{syn}}\cdot n_{2}^{\text{syn}}}\\ \bm{1}^{T}\bm{b}=m^{\text{syn}}\end{subarray}}\lVert\frac{1}{m^{\text{syn}}}\hat{\bm{Q}}\bm{b}-\hat{\bm{a}}\rVert_{2}^{2}$$\triangleright$ Section 4
    $\bm{B}^{\text{syn}}=\operatorname{UnbiasedSampling}(\widetilde{\bm{b}}^{\text{syn}},m^{\text{syn}})$$\triangleright$ Section 5